# Channel Choice and Customer Value

### v1

@@ 5 Firm’s Problem and Sample Construction
We first discuss the sample construction for the three focal events in $\S$5.1 and then present some summary statistics of the three samples in $\S$5.2.
@@ 5.1 Sample Construction
For each event, we first define three customer cohorts – event adopters, organic adopters, and offline-only customers. We then establish the pre-adoption, adoption, and post-adoption periods for each focal event. Specifically, for a given event:
@@
See Table 5 for an overview of these periods for each of the three events. In our estimation sample for each event, we only include data from the pre- and post-adoption periods for that event. We exclude the adoption periods for both event and organic adopters from the estimation sample because consumers are in flux during that time (where some adopters have adopted while others have not). Furthermore, by excluding these observations, we avoid capturing any abnormal purchase behaviors associated with the initial adoption of online shopping and the onset of specific events (i.e., COVID-19, Black Friday, Loyalty Program). As such, it allows us to cleanly identify the post-online-adoption difference in customers’ purchase behaviors (Gu and Kannan 2021). We now clarify a few points on the choice of these three periods for each event below.
@@
    COVID-19 event: We use one year of data before the adoption period as the pre-adoption period (the calendar year of 2019). For post-adoption period, we use all the data from May 2020 till April 2023666Pilot studies for the firm’s Loyalty program started running from May 2023; so we cut off the post-adoption period with April 2023..
@@ 5.1.1 Event 1: COVID-19
The COVID-19 pandemic was an exogenous macro-environmental shock that prompted many consumers to adopt online shopping. Figure 1(a) shows the number of customers who adopted online shopping through the retailer’s website and/or mobile app (from previously only shopping at physical stores) for each month for our observation period. As we can see, during March and April 2020, there was a large peak in the number of offline customers who first switched to online shopping; the number of adopters sharply increased from approx. 7,000 in January and February 2020 to approximately 20,000 in March and April 2020777In Figure 1(a), we also see smaller peaks associated with the second and third waves of COVID-19.. The customers who adopted online shopping during the onset of COVID-19 may have done so for reasons like health concerns, social distancing norms, etc., whereas the customers who adopted online shopping before that did so of their own volition, for other reasons (e.g., the convenience of online shopping), unrelated to COVID-19.
@@
    Covid adopters: Customers who made their first online purchase during the onset of COVID-19, i.e., from 2020-03-16 8882020-03-16 was the date when the Brazilian government started implementing physical distancing and confinement measures in response to COVID-19 (Faria de Moura Villela et al. 2021). to 2020-04-30.  [^9] The set of people who adopted during this period could also include consumers who would have organically adopted even without the COVID-19 shock. We have no clear way of distinguishing such potential organic adopters from those who adopted online shopping as a result of COVID-19. Nevertheless, based on Figure 1(a), it seems like approximately two-thirds of these consumers switched due to COVID-19. That is, a significant majority of these consumers switched due to COVID-19, and we refer to this cohort as covid adopters for simplicity. Please see $\S$8.1 for a detailed discussion on the implications of this potential mixing on the ATT estimates for covid and the other two events.
@@ 5.1.2 Event 2: Black Friday Sales
For this estimation sample, we consider the cohort of customers who made their first offline purchase before 2019-10-31 and continued to be the firm’s customers after the Black Friday sales date. Then, we categorize each consumer from this cohort into one of the following three groups (see Table third column in 5 for exact dates):
@@ 5.1.3 Event 3: Loyalty Program
Finally, we consider a retailer-led marketing campaign that incentivized offline-only customers to adopt the online channel. In August and September 2023111111Pilot studies for the loyalty program ran in two states (Rio Grande do Sul and Paraná) as early as May. Therefore, we exclude these two states from the estimation sample for this event., the retailer launched a loyalty program and extensively advertised it to customers through email and store marketing. Customers who wanted to join the LP had to use the mobile app (no offline options were available for activating the loyalty program). After completing the online activation, customers could access and activate monthly special offers and receive discounts on both online and offline purchases [^12] Note that customers had to activate each month for participation in each month’s special offers.. Figure 1(c) presents the number of adopters by month since January 2023 and shows clear peaks in the number of adopters during August and September 2023, coinciding with the launch of the loyalty program. This suggests that online-only activation of the newly available loyalty program prompted a significant number of customers to adopt online shopping during the roll-out period.
For this event, we consider the cohort of customers who made their first offline purchase before 2022-12-31 and continued to be the firm’s customers during the post-online-adoption period (i.e., after 2023-10-01). Then, we categorize each consumer from this cohort into one of the following groups (see last column of Table 5 for details):
@@ 5.2 Summary Statistics
For each event (Covid, Black Friday, and Loyalty Program), we generate customer-month-level panel data for the pre- and post-adoption periods. These three panel datasets will form the basis of our empirical analysis going forward. We present the detailed summary statistics on demographics (gender, age, household income, and locations), and pre-adoption purchase behaviors (spend, quantity, orders, brands, categories, etc.) for the three groups of customers (offline-only, organic adopters and event adopters) from each event (Covid, Black Friday and Loyalty Program) in Web Appendix $\S$§C. We discuss some key takeaways here.
@@
However, there are some important differences between the two groups of adopters (event adopters and organic adopters) across all three events. For instance, in the case of COVID-19, event/covid adopters tend to be older, have higher incomes, and have been shopping with the retailer for a longer time compared to organic adopters. However, these two groups have similar pre-adoption monthly spending. In the case of Black Friday, event(or Black Friday) adopters are younger, have been shopping with the retailer for a longer time, and have lower monthly spend with the retailer in the pre-adoption period relative to organic adopters. In contrast, in the Loyalty program case, event adopters tend to be slightly older, have been shopping with the retailer for a shorter period, and have a higher monthly spend with the retailer in the pre-adoption period compared to organic adopters. However, the differences in characteristics between the two groups of adopters for each event are much smaller in magnitude relative to the differences between adopters and non-adopters (i.e. offline-only customers).
6 Empirical Specification and Results
We now present our DID models in $\S$6.1 and the results in $\S$6.2.
6.1 Model
To estimate the impact of adopting online shopping on post-adoption purchase behavior, we use the difference-in-differences approach. Unlike our descriptive analyses, this approach allows us to account for unobserved heterogeneity across customers and time-specific shocks. For each event, we use customer-month-level panel data for the estimation sample corresponding to that event, and we utilize the customer-month observations from the pre- and post-adoption periods. We exclude the adoption periods from the estimation for the reasons discussed in $\S$5.1.
@@ 6.1 Model
We have three main outcome variables of interest: spend, profitability and share of offline spend. We now discuss the interpretation of the coefficients and variables in the above equation for the case when spend is the outcome variable of interest. The interpretations are similar for the other outcome variables.
The intercept $\alpha_{0}$ captures the average monthly spend of offline-only customers in the pre-adoption period. $\mathbbm{I}\{Adopter_{i}\}$ is a binary variable that takes the value 1 for both organic adopters and event adopters, and 0 for offline-only customers. $\alpha_{1}$ denotes the pre-adoption period spend of adopters (including both organic adopters and event adopters), compared to offline-only customers. $\mathbbm{I}\{Post_{t}\}$ is a binary variable that takes the value 1 for the post-adoption periods, and 0 for pre-adoption periods. $\alpha_{2}$ captures the average change in spend of offline-only customers in the post-adoption period, relative to the pre-adoption period. $\alpha_{3}$ represents the average change in spend per month for adopters in the post-adoption period, respectively, relative to offline-only customers. $\alpha_{3}$ is the coefficient of interest here, representing the ATT estimate. We also include other time-varying control variables $\boldsymbol{{x}}_{it}$ to better isolate the causal effect of online adoption from other confounding factors that affect the outcome variables for one of our events – the loyalty program. Specifically, we construct control variables that capture customers’ participation in the loyalty program, such as whether they activate the offer or not [^14] We describe the list of control variables in Web Appendix $\S$§C.4..
@@
Here, $(\delta_{3})$, the coefficient of interest, represents the difference in the change in spend between event adopters and organic adopters in the post-adoption period relative to pre-adoption. We also include a second specification, Equation (4) with customer fixed effects ($\gamma_{i}$) and time (month) fixed effects ($\eta_{t}$).
@@
The validity of all the DID models discussed here relies on the parallel trends assumption. We discuss the validity of this assumption for all the models and outcome variables in $\S$8.2.
6.2 Results
We estimate the DID models specified above for two of our main outcome variables of interest: spend and profit, and present the results in this section. We discuss our third variable of interest, share of offline/online spend, in the next section. We also consider other outcome variables, e.g., quantity, orders, and number of unique brands purchased, in Web Appendix $\S$§E.2.
6.2.1 Effect on Total Spend
First, we present the results from the DID models specified in Equations (1) and (2) with offline-only customers as the control group and spend as the outcome variable in Table 6. Across all three events (COVID-19, Black Friday, and Loyalty Program), we find that adopters (pooling organic and event adopters) increase their spending in the post-adoption period (relative to pre-adoption) compared to offline-only customers.
Next, to examine the differences between each of the different event adopters and organic adopters, we estimate the DID models specified in Equations (3) and (4). The results from Equation (3) are in column (1), (3), and (5) of Table 7 for Covid, Black Friday, and Loyalty Program, respectively. We see that there is no significant difference in the post-adoption spend between covid adopters and organic adopters. However, the incremental spending of Black Friday adopters and Loyalty Program adopters is lower than organic adopters, post-adoption. We also control for customer and time fixed effects as shown in Equation (4) (in column (2), (4), and (6) of Table 7) and confirm the robustness of these findings to these additional controls.
In summary, our findings suggest that the different reasons for the adoption of online shopping do matter for the post-adoption incremental spending. While adopters of online shopping increased their spending with the retailer in the post-adoption period, relative to offline-only customers; those who adopt due to a macro-environmental shock (COVID-19, in our case) spend similarly post-adoption, relative to organic adopters, and those who adopt due to a promotional activity (Black Friday and Loyalty Program) by the firm, spend less.
Table 6: DID Main Analysis - Spend (Adopters vs. Offline-only)
@@ 6.2.1 Effect on Total Spend
Table 7: DID Main Analysis - Spend (Event Adopters vs. Organic Adopters)
@@
6.2.2 Effect on Profitability
@@
In this section, we examine whether and how the adoption of online shopping affects the profitability of adopters relative to offline-only customers, and whether there are differences between the different groups of event adopters relative to organic adopters.
@@
In our data, we observe item-level profit margins for each purchase [^15] We observe this starting January 2020. We obtain the 2019 profit margins of each item in a given channel by extrapolating the item-channel level profit margins from 2020–2023 to the 2019 time-frame based on Brazil’s Extended National Consumer Price Index (referred to as IPCA). We refer interested readers to Web Appendix $\S$§D for details. From our discussions with managers at the focal retailer, we do not have any reason to believe that the profit margin structures changed between 2019 and the rest of the data period.. We first provide some summary statistics of the differences in prices and profit margins across the two channels. To highlight the differences in these three variables across channels, we first generate a panel of prices at the item-channel-month level [^16] For instance, in the case of price, an observation in this panel is the average price of an item $j$ sold through a channel $c$ in time period $t$, and is calculated by the total sales of item $j$ in channel $c$ in month $t$ divided by the total quantity of that item sold in that channel-month.. Next, we calculate the offline-online differences at the item-month level to obtain item-specific offline-online differences in price and profit margin. We then average over months and obtain item-level average differences in price and profit margin between offline and online channel. Table 8 provides item-level summary statistics for the 22,803 items that were sold in both offline and online channels in our data (and were purchased at least once in the same month). For confidentiality, we multiply the profit margin by the same undisclosed multiplier used for the price. We see from Table 8, that on average, prices in physical stores are higher than online prices, and profit margins are higher in physical stores vs. online. Thus, consumers who spend more in physical stores may be more profitable (assuming they buy similar assortments of products).
@@
Table 9: DID Main Analysis - Profit (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV: Profit
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 92.57 (0.177) 96.13 (0.187) 125.8 (0.288)
[0.000] [0.000] [0.000]
Adopter 61.35 (0.925) 60.94 (3.041) 41.57 (1.607)
[0.000] [0.000] [0.000]
Post -10.56 (0.174) -22.83 (0.152) -16.81 (0.238)
[0.000] [0.000] [0.000]
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
R2 0.01228 0.31903 0.00450 0.34750 0.00651 0.44546
Within R2 0.00047 $2.48\times 10^{-5}$ 0.00193
Next, we estimate the DID models in Equations (1)–(2) with profit as the dependent variable, and present the results in Table 9. We see that across all the events, adopters show an increase in profitability post-adoption, relative to offline-only customers.
@@
Finally, we investigate the differences between each group of event adopters relative to organic adopters by running the DID models specified in Equation (3) and Equation (4). The results are shown in Table 10. First, we see that covid adopters are significantly more profitable than organic adopters post-adoption. The differences are economically meaningful and large in magnitude – the post-adoption gain in profitability is over 30% (7.248/20.27) higher for covid adopters relative to organic adopters. The fact that covid adopters are more profitable than organic adopters despite there being no significant differences in spend is an interesting finding. This may be due to the differences in prices across channels in our setting and potential differences in the share of offline/online spending. We explore this mechanism more in the following section. Also interestingly, we find that Black Friday adopters and Loyalty Program adopters are less profitable relative to organic adopters in the post-adoption period by 10.2% (16.04 / (154.7 + 3.3)) and 10.7% (15.41 / (164.5 - 21.05)), respectively, and this is likely driven by their lower spending post-adoption.

### v2

@@ 5 Estimation Sample and Summary Statistics
We first discuss the sample construction process for the three focal events in $\S$5.1, and present the details of the samples for each of the three focal events in $\S$5.2, and then present some summary statistics of the three samples of interest in $\S$5.3.
@@ 5.1 Sample Construction
We now discuss how we construct the samples used in our empirical analyses and the main outcome variables of interest.
Focal events that triggered different pathways of adoption: Since we seek to connect channel adoption pathway to post-adoption behavior, we consider three events that each led to a surge in online adoption: (1) a macro-environmental shock (COVID-19), (2) a short-term promotion (Black Friday sales), and (3) a long-term loyalty program. The first event was a macroeconomic shock outside the firm’s control, whereas the latter two were firm-driven marketing activities.
Customer cohorts: For each event, we first define three customer cohorts – (a) event adopters, consumers who adopted online shopping due to the event, (b) organic adopters, consumers who adopted organically (without external incentives) around the same time frame as the event, and (c) offline-only customers, consumers, who did not adopt online shopping at all.
Time periods: Next, we establish the pre-adoption, adoption, and post-adoption periods for each focal event as follows:
@@
See Table 2 for an overview of these periods for each of the three events. In our estimation sample for each event, we only include data from the pre- and post-adoption periods for that event. We exclude the adoption periods for both event and organic adopters from the estimation sample because consumers are in flux during that time (where some adopters have adopted while others have not). Furthermore, by excluding these observations, we avoid capturing any abnormal purchase behaviors associated with the initial adoption of online shopping and the onset of specific events (i.e., COVID-19, Black Friday, Loyalty Program). As such, it allows us to cleanly identify the post-online-adoption difference in customers’ purchase behaviors; see Gu and Kannan (2021) for a similar specification. We now summarize the choice of these three periods for each event.
@@
    COVID-19 event: We use one year of data before the adoption period as the pre-adoption period (the calendar year of 2019). For the post-adoption period, we use all the data from May 2020 till April 2023444Pilot studies for the firm’s Loyalty program started running from May 2023; so we cut off the post-adoption period with April 2023..
@@
Outcome metrics: For all our analyses, we consider three metrics that quantify a consumer’s shopping behavior in a given month: (1) Total Spend, (2) Profitability, and (3) Stickiness to the offline channels post-adoption. Formally, we are interested in the Average Treatment Effect on Treated (ATT) of adopting online shopping on these three metrics for each of the three focal events. Each of these metrics is defined at the customer-month level.
5.2 Estimation Samples for the Focal Events
We now describe the estimation sample used in the analysis of each of the three focal events.
@@ 5.2.1 Event 1: COVID-19
The COVID-19 pandemic was an exogenous macro-environmental shock that prompted many consumers to adopt online shopping. Figure 1(a) shows the number of customers who adopted online shopping through the retailer’s website and/or mobile app (from previously only shopping at physical stores) for each month for our observation period. As we can see, during March and April 2020, there was a large peak in the number of offline customers who first switched to online shopping; the number of adopters sharply increased from approx. 7,000 in January and February 2020 to approximately 20,000 in March and April 2020555In Figure 1(a), we also see smaller peaks associated with the second and third waves of COVID-19.. Customers who adopted online shopping during the onset of COVID-19 may have done so for reasons like health concerns, social distancing norms, etc., whereas the customers who adopted online shopping before that did so of their own volition, for other reasons (e.g., the convenience of online shopping), unrelated to COVID-19.
@@
    Covid adopters: Customers who made their first online purchase during the onset of COVID-19, i.e., from 2020-03-16 6662020-03-16 was the date when the Brazilian government started implementing physical distancing and confinement measures in response to COVID-19 (Faria de Moura Villela et al. 2021). to 2020-04-30.  [^7] The set of people who adopted during this period could also include consumers who would have organically adopted even without the COVID-19 shock. We have no clear way of distinguishing such potential organic adopters from those who adopted online shopping as a result of COVID-19. Nevertheless, based on Figure 1(a), it seems like approximately two-thirds of these consumers switched due to COVID-19. That is, a significant majority of these consumers switched due to COVID-19, and we refer to this cohort as covid adopters for simplicity. Please see $\S$9.1 for a detailed discussion on the implications of this potential mixing on the ATT estimates for COVID-19 and the other two events.
@@ 5.2.2 Event 2: Black Friday Sales
For this estimation sample, we consider the cohort of customers who made their first offline purchase before 2019-10-31 and continued to be the firm’s customers after the Black Friday sales date. Then, we categorize each consumer from this cohort into one of the following three groups (see third column in Table 2 for exact dates):
@@ 5.2.3 Event 3: Loyalty Program
In August and September 2023999Pilot studies for the loyalty program ran in two states (Rio Grande do Sul and Paraná) as early as May. Therefore, we exclude these two states from the estimation sample for this event., the retailer launched a loyalty program and extensively advertised it to customers through email and store marketing. Customers who wanted to join the LP had to use the mobile app (no offline options were available for activating the loyalty program). After completing the online activation, customers could access and activate monthly special offers and receive discounts on both online and offline purchases [^10] Note that customers had to activate each month for participation in each month’s special offers.. Figure 1(c) presents the number of adopters by month since January 2023 and shows clear peaks in the number of adopters during August and September 2023, coinciding with the launch of the loyalty program. This suggests that online-only activation of the newly available loyalty program prompted a significant number of customers to adopt online shopping during the roll-out period. For this event, we consider the cohort of customers who made their first offline purchase before 2022-12-31 and continued to be the firm’s customers during the post-online-adoption period (i.e., after 2023-10-01). Then, we categorize each consumer from this cohort into one of the following groups (see last column of Table 2 for details):
@@ 5.3 Summary Statistics
For each consumer in the estimation sample of any focal event, we aggregate their transaction-level data into customer-month-level panel data over the customer’s entire purchase history within the relevant period (January 2019 to March 2024), starting from their initial purchase month and ending with the last observed purchase month. For example, if a customer purchased for the first time in our data in August 2019 and the last observed purchase is in July 2020, then the panel for this customer spans 12 months, from August 2019 to July 2020. If there are intermediate months in which a customer did not make any purchase, the values for the monthly purchase-related variables are set to zero.
[…]
@@
However, there are some important differences between the two groups of adopters (event adopters and organic adopters) across all three events. For instance, in the case of COVID-19, event/covid adopters tend to be older, have higher incomes, and have been shopping with the retailer for a longer time compared to organic adopters. However, these two groups have similar pre-adoption monthly spending. In the case of Black Friday, event/Black Fridayadopters are younger, have been shopping with the retailer for a longer time, and have lower monthly spend with the retailer in the pre-adoption period relative to organic adopters. In contrast, in the Loyalty program case, event adopters tend to be slightly older, have been shopping with the retailer for a shorter period, and have a higher monthly spend with the retailer in the pre-adoption period compared to organic adopters. However, the differences in characteristics between the two groups of adopters for each event are much smaller in magnitude relative to the differences between adopters and non-adopters (i.e. offline-only customers).
6 Empirical Strategy
As discussed earlier, our goal is to estimate the ATT (Average Treatment Effect on the Treated) of online adoption on post-adoption purchase behavior for different types of events that trigger the adoption of online shopping, and then connect these ATTs estimates with the different adoption pathways. To estimate our ATTs of interest, we use the difference-in-differences approach. This approach allows us to account for unobserved heterogeneity across customers and time-specific shocks. For each event, we use customer-month-level panel data for the estimation sample corresponding to that event, and we utilize the customer-month observations from the pre- and post-adoption periods. We exclude the adoption periods from the estimation for the reasons discussed in $\S$5.1.
@@ 6 Empirical Strategy
As discussed earlier, we have three main outcome variables of interest: spend, profitability, and share of offline spend. We now discuss the interpretation of the coefficients and variables in the above equation for the case when spend is the outcome variable of interest. The interpretations are similar for the other outcome variables.
[…]
@@
$(\delta_{3})$, the coefficient of interest, represents the difference in the change in spend between event adopters and organic adopters in the post-adoption period relative to pre-adoption. We also include a specification, Equation (4) with customer fixed effects ($\gamma_{i}$) and time (month) fixed effects ($\eta_{t}$).
@@
The validity of all the DID models discussed here relies on the parallel trends assumption, which we discuss in detail in $\S$9.2. Further, we discuss in $\S$9.1 how potential contamination between the different types of adopters could affect our results and conclusions. For instance, some adopters who we identify as covid adopters because they adopt during March-April 2020 could actually be organic adopters who would have adopted organically during that time anyway. If this is the case, we show that our measured effect would be an underestimate of the true the difference between event adopters and organic adopters.
7 Results
7.1 Adopters vs. offline-only customers
[…]
The results from the DID models specified in Equations (1) and (2) with offline-only customers as the control group and spend as the outcome variable are presented in Table 3. Across all three events (COVID-19, Black Friday, and Loyalty Program), we find that adopters (pooling organic and event adopters) increase their spending in the post-adoption period (relative to pre-adoption) compared to offline-only customers. To measure effect sizes on a relative scale, we report each ATT as a percentage of pre-adoption monthly spend. The post-adoption monthly spend increases by 30.6% for COVID-19 (i.e., 140.1 / (280.1 + 177.2)), 16.7% for Black Friday (i.e., 78.37 / (291.7 + 178.2)), and 16.6% for the Loyalty Program (i.e., 84.89 / (332.6 + 179.2)). It is interesting to note that the macro-shock (COVID-19) and promotions (Black Friday and Loyalty Program) yield differential increases in spending. We explore this more in the following sub-section.
Table 3: DID Main Analysis – Spend (Adopters vs. Offline-only)
@@ 7 Results
[…]
Table 4: DID Main Analysis - Profit (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV: Profit
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 92.57 (0.177) 96.13 (0.187) 125.8 (0.288)
[0.000] [0.000] [0.000]
Adopter 61.35 (0.925) 60.94 (3.041) 41.57 (1.607)
[0.000] [0.000] [0.000]
Post -10.56 (0.174) -22.83 (0.152) -16.81 (0.238)
[0.000] [0.000] [0.000]
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
R2 0.01228 0.31903 0.00450 0.34750 0.00651 0.44546
Within R2 0.00047 $2.48\times 10^{-5}$ 0.00193
In summary, we see that consumers who adopt online shopping both spend more and are more profitable than offline-only users, across all three events, confirming our prediction H1. This finding is consistent with the earlier literature on multichannel shopping (Neslin and Shankar 2009; Montaguti, Neslin, and Valentini 2016). However, as discussed earlier, both prior work and this analysis treat all adopters as a homogeneous unit, i.e., they ignore channel adoption pathways. As such, it masks any heterogeneity in how channel pathways affect post-adoption spend, profitability, and purchase behavior. In the next section, we present empirical findings on the relationship between channel adoption pathways and post-adoption consumer behavior.
7.2 Event adopters vs. organic adopters
We now examine the differences between each of the different event adopters and organic adopters by estimating the DID models specified in Equations (3) and (4). We will first present the results on spend, followed by the share of offline/online spending, and then look at profitability. These results will help us answer our second research question, and verify the second set of predictions (H2a through H2d) presented in $\S$3.
7.2.1 Spend
The results from the estimation of Equation (3) for spend are in columns (1), (3), and (5) of Table 5 for Covid, Black Friday, and Loyalty Program, respectively. We see that there is no significant difference in the post-adoption spend between covid adopters and organic adopters. However, the incremental spending of Black Friday adopters and Loyalty Program adopters is lower than organic adopters, post-adoption. We also control for customer and time fixed effects as shown in Equation (4) (in columns (2), (4), and (6) of Table 5) and confirm the robustness of these findings to these additional controls.
Table 5: DID Main Analysis – Spend (Event Adopters vs. Organic Adopters)
@@ 7.2.1 Spend
These findings suggest that the reason why a consumer adopts online shopping affects the post-adoption incremental spending. While all adopters of online shopping increase their spending with the retailer in the post-adoption period, relative to offline-only customers, those who adopt due to a macro-environmental shock (COVID-19, in our case) spend similarly post-adoption, relative to organic adopters. In contrast, those who adopt due to a promotional activity (Black Friday and Loyalty Program) by the firm, spend less than organic adopters. This is an important finding for managers and researchers since it suggests that ignoring channel adoption pathways can significantly misestimate the incremental gains from online adopters. Further, it is consistent with our theoretical prediction (H2a) of forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Hendel and Nevo 2006; Van Heerde and Neslin 2017) from $\S$3. We know that promotions are an important mechanism for firms to induce various kinds of consumer behavior, including the adoption of online shopping. However, time-limited promotions may induce forward buying, leading to demand distortion and reduced post-promotion consumption.
Theory and Mechanism: Black Friday adopters and Loyalty Program adopters could be increasing their spending during the promotion (i.e., the adoption period) and effectively pulling forward their future demand, resulting in a “post-promotion dip” in spending in the post-adoption period.
To formally examine if this is the case, we investigate the differences between the two groups of event adopters relative to organic adopters during each month around adoption. We estimate the following event study regression:
[…]
(a) Black Friday
(b) Loyalty Program
7.2.2 Share of offline/online spend
As discussed in $\S$3, consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009)would predict that adoption driven by external shocks (e.g., COVID-19, firm driven promotions) may lead to temporary disruptions in consumer habits (adopting a new channel and multichannel shopping), but without reinforcement, consumers may revert to prior routines (not use the newly adopted channels much). We test this prediction (H2b) in our empirical setting, by comparing post-adoption channel usage between consumers who adopted online shopping organically and those who did so in response to the COVID-19 shock, a short-term one-time firm-driven promotion (Black Friday) and a more long-term and recurring firm driven promotion (Loyalty Program). In other words, we examine if there are differences in how the total spend with the retailer is distributed across the offline and online channels for the different groups of event adopters relative to organic adopters in the post-adoption period.
[…]
(a) Covid
(b) Black Friday
(c) Loyalty Program
[…]
7.2.3 Profitability
@@
Now, we examine the differences in profitability between the different groups of event adopters relative to organic adopters. We know that there are differences in spend and in the share of offline/online spending.
@@
There are also differences in prices and profit margins across the two channels in our setting. To highlight the differences in these three variables across channels, we first generate a panel of prices at the item-channel-month level [^21] For instance, in the case of price, an observation in this panel is the average price of an item $j$ sold through a channel $c$ in time period $t$, and is calculated by the total sales of item $j$ in channel $c$ in month $t$ divided by the total quantity of that item sold in that channel-month.. Next, we calculate the offline-online differences at the item-month level to obtain item-specific offline-online differences in price and profit margin. We then average over months and obtain item-level average differences in price and profit margin between offline and online channel. Table 6 provides item-level summary statistics for the 22,803 items that were sold in both offline and online channels in our data (and were purchased at least once in the same month). For confidentiality, we multiply the profit margin by the same undisclosed multiplier used for the price.
@@
Thus, on average, prices in physical stores are higher than online prices, and profit margins are higher in physical stores vs. online. Therefore, even for similar levels of spend, consumers who have a higher share of offline spend may be more profitable (our prediction H2d). We present the results from estimating the DID models specified in Equation (3) and Equation (4) for profitability are shown in Table 7.
@@
We see that covid adopters are significantly more profitable than organic adopters post-adoption. The differences are economically meaningful and large in magnitude – the post-adoption gain in profitability is over 30% (i.e., 7.248/20.27) higher for covid adopters relative to organic adopters, consistent with our prediction H2d. The fact that covid adopters are more profitable than organic adopters despite there being no significant differences in spend is due to the differences in prices across channels in our setting and the differences in the share of offline/online spending, as documented in $\S$7.2.2. This shows that the differences in share of spend across channels arising due to consumer habit and inertia have significant economic implications for firms. Also, we find that Black Friday adopters and Loyalty Program adopters are less profitable relative to organic adopters in the post-adoption period by 10.2% (i.e., 16.04 / (154.7 + 3.3)) and 10.7% (i.e., 15.41 / (164.5 - 21.05)), respectively, and this is driven by their lower spending post-adoption as shown in $\S$7.2.1.